OpenStudy (anonymous):

factor a63+27= (a-3)(a^2+3a+9)?

OpenStudy (anonymous):

oops its a^3+27

OpenStudy (anonymous):

rather (a+3)(a^2-3a+9)

OpenStudy (anonymous):

aah. thanks.

OpenStudy (anonymous):

glad i asked

OpenStudy (paxpolaris):

\[\huge a^3+b^3=\left( a+b \right)\left( a^2-ab+b^2 \right)\]